The half-life of tritium is 12.4 years. How long will it take for 15% of a sample of tritium to decompose?
An isotope of thorium, , has a half-life of 18.4 days. How long will it take for 77% of the sample to decompose?






Answer by stanbon(75887)

The half-life of tritium is 12.4 years. How long will it take for 15% of a sample of tritium to decompose?
If 15% decomposes, 85% remains.
-----------------
A(t) = Ao*e^(-ln(2t)/12.4)
0.85 = e^(-ln(2t)/12.4)
0.85 = 2t/12.4
2t = 10.54
t = 5.27 years


An isotope of thorium, , has a half-life of 18.4 days. How long will it take for 77% of the sample to decompose?
---
0.23 = e^(-ln(2t)/18.4)
0.23 = 2t/18.4
2t = 4.232
t = 2.116 days
=======================
Cheers,
Stan H.